How many grams of FeCl3 are present in 15.00 mL of the solution? (Molar mass of FeCl3 = 162.2 g/mol)


1
Expert's answer
2022-03-12T01:58:04-0500

Solution:

"V=15.00mL=15.00\sdot10^{-3}L;"

"M(FeCl_3)=162.2 g\/mol;"

"\nu=\dfrac{V}{V_{M}};"

"\nu=\dfrac{15.00\sdot10^{-3}}{22.4}=0.67\sdot 10^{-3}mol;"

"\nu=\dfrac{m}{M(FeCl_3)}\implies m=\nu\sdot M(FeCl_3);"

"m=0.67\sdot10^{-3}mol\sdot162.2g/mol=108.67\sdot 10^{-3}g=0.109g;"

Answer: "m=0.109g."